40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 31 07:41:23 2011, Aidan, Update, Green Locking, DFD - medm screen Screenshot-3.png
    Reply  Mon Jan 31 12:40:38 2011, rana, joe, Update, Green Locking, DFD - medm screen a.pdf
       Reply  Mon Jan 31 18:25:25 2011, Aidan, Update, Green Locking, DFD - results from the new filters (and running with AWG) DFD-bandwidth_noise_CLP500_to_300Hz.pdfDFD-bandwidth_noise_CLP500_to_7000kHz.pdf
Message ID: 4234     Entry time: Mon Jan 31 18:25:25 2011     In reply to: 4232
Author: Aidan 
Type: Update 
Category: Green Locking 
Subject: DFD - results from the new filters (and running with AWG) 

Quote:

This is a plot showing the old filters and the new ones we added this morning.

The new ones have a Cheby for AC coupling below 10 Hz and then a 500 Hz LP after the mixer. The LP frequency has been increased so that we can use this signal in a feedback loop to the ETM with a ~100 Hz UGF.

Joe injected a 234.567 etc. Hz sine wave into the excitation channel in the DFD INPUT filter. The spectrum of the output of the LP filter with the new filter is shown below with the RMS calculated from 300Hz down to 1mHz - see first attachment. The RMS is equal to about 2.5Hz. (Incidentally, the RMS is very much higher (slightly less than 400Hz - see second attachment) if you calculate it from 7kHz down to 1mHz). 

Attachment 1: DFD-bandwidth_noise_CLP500_to_300Hz.pdf  22 kB  | Hide | Hide all
DFD-bandwidth_noise_CLP500_to_300Hz.pdf
Attachment 2: DFD-bandwidth_noise_CLP500_to_7000kHz.pdf  15 kB  | Hide | Hide all
DFD-bandwidth_noise_CLP500_to_7000kHz.pdf
ELOG V3.1.3-